# AutoCodeRover: Autonomous Program Improvement

### v2

@@ 6.2. RQ2: Effect of SBFL
In this research question, we want to understand whether program analysis techniques such as Spectrum-based fault localization (SBFL) can benefit the workflow of AutoCodeRover. Different from RQ1, here we construct a common scenario in program repair that AutoCodeRover has access to the complete test-suite of the target task instance. We use the developer-written test cases for each task instance (provided in SWE-bench lite) as the test-suite. To understand the effect of SBFL, we perform two sets of experiments: (1) AutoCodeRover and use the test-suite for patch validation during the patch generation retry-loop; (2) additionally provide SBFL results (top-5 suspicious methods) to AutoCodeRover at the beginning of the context retrieval stage compared to (1). We denote these two settings as ACR-val and ACR-val-sbfl. The patch validation is added as follows: when a patch is generated by the LLM agent, the test-suite is executed on the patched program; if the patch fails to pass the complete test-suite, AutoCodeRover re-invokes the patch generation agent to write a new patch. This validation loop is configured to run at most three times.
Table 4. Result of ACR-val, ACR-val-sbfl on SWE-bench lite, in one run only.
Tools Resolved Tasks Avg Time Avg Tokens
ACR-avg 16.11% (48.3) 173 37602 (\$0.435)
ACR-val 17.00% (51) 234 46629 (\$0.538)
ACR-val-sbfl 20.33% (61) 250 40545 (\$0.470)
@@ Results.
Table 4 shows that, with the additional information provided by SBFL, the number of resolved tasks increased from 51 to 61 (i.e. 17.00% to 20.33% resolved rate on SWE-bench lite). Compared to the validation-only setting, additionally adding the SBFL component helps to resolve 20 unique task instances. Moreover, when comparing with task instances resolved in all other runs combined (i.e. ACR-all and ACR-val) on SWE-bench lite, ACR-val-sbfl still uniquely resolves 7 task instances that are not resolved in any of the other runs.
@@
We present a case study on one of the tasks uniquely resolved by ACR-val-sbfl. Figure 8 shows the issue description of django-13964555https://code.djangoproject.com/ticket/32332 and how AutoCodeRover attempts to retrieve code context with and without SBFL. This issue reported a bug when saving django models to a database. A simplified version of the issue description is shown in the first part of Figure 8, in which important parts are highlighted. In this issue, some “hints” (highlighted in red) mentioned are actually distracting factors for a context retrieval agent. For example, the Product and Order classes describe how the bug can be reproduced, and are not classes that cause the bug. Retrieving code context from these two classes (as shown in Figure 8 part 2 “WITHOUT SBFL”) does not yield useful results for forming the context and resolving the issue.
@@ 6.3. RQ3: Challenges on real-life tasks
In this research question, we analyze the task instances in SWE-bench lite that AutoCodeRover failed to resolve, and provide a taxonomy of the issue characteristics to highlight the practical challenges in achieving fully automated software improvement. Our taxonomy consists of challenges in the fault localization stage and patch generation stage. Specifically, for each task, we analyze the best run in the three repetitions, and classify each of the 300 tasks into one of the following:
@@
Figure 9 shows the distribution of the 300 tasks in SWE-bench lite. AutoCodeRover resolves 22.3% of the issues (“Success”), as mentioned in Section 6.1. The fail-to-resolve cases are included in the remaining four categories. In 31.7% of the tasks, AutoCodeRover correctly decided on all patch locations (at the method-level), but did not produce a correct patch (“Wrong patch”). More fine-grained intra-procedural analysis and specification inference techniques can play a significant role in improving these cases, by providing the patch generation agent with more method-level repair guidance. In the other three categories, the fault localization could not pinpoint all the locations to be modified. In 20.7% of the tasks, a patch is generated in the correct file, but at wrong methods / classes in the file (“Wrong location in correct file”). In some of these runs, the developer patch modifies multiple methods, but the generated patch did not modify all of them. In the other categories, a patch could not be generated at the correct file - in 18.7% of the tasks a patch is generated in wrong files, and in 6.7% of the tasks there is no applicable patch (“Wrong file” and “No patch”). We manually inspected some tasks in these two categories, and observed that their issue description mentions few methods / classes / files in the codebase. Instead, some of them contain short examples to reproduce the issue. For these tasks, one possibility is to generate a comprehensive test-suite based on the issue description, and then use execution information of the test-suite (e.g. SBFL) to reveal suspicious program locations. On the other hand, some other tasks do not contain reproducible examples and only consist of natural language descriptions. For these tasks some human involvement might be helpful. The developers could focus on these tasks.
7. Discussion on improvements
In this section, we discuss a few possible directions for future improvements on AutoCodeRover.
@@ 7. Discussion on improvements
In Section 4, we described two scenarios that AutoCodeRover can work on, (1) issue description only, (2) aided by SBFL when the test suite is available. Although the second scenario is not always practical in real-life software development, the issue description sometimes contains a concrete reproduction script from the user that reports the bug. In the future, it is possible to design an LLM agent specifically to generate a bug reproduction test based on the GitHub issue description. The bug reproduction test can then be used to validate the correctness of AutoCodeRover’s generated patch and possibly enable a regeneration process if a patch fails to pass the reproduction test. However, we also note that it is not an easy task, because large projects often use different automated testing frameworks, a test that cannot be properly executed may even confuse the main LLM agent on the GitHub issue and lead to a negative result.
@@
During context retrieval, AutoCodeRover navigates through the codebase by visiting code entities such as classes and methods, which has shown to be effective. The idea of utilizing code-specific structure for context retrieval can be taken further by considering artifacts from program semantics. For example, from an initial set of methods identified in the issue description, a static call graph analysis (Ryder, 1979; Salis et al., 2021) can be used to collect additional relevant methods when testcases are absent. There are also potential in integrating a language server (Gunasinghe and Marcus, 2021) for codebase navigation, so that the context retrieval agent can perform more code-specific actions such as “jump-to-definition” from a method invocation. Finally, one can use forward data dependence analysis (Jin et al., 2024) from the methods in the issue description to search for other relevant methods.
@@
Currently, AutoCodeRover left all decision-making processes to its foundation LLM — e.g. deciding context retrieving locations and whether to terminate the retrieval stage. However, leaving it to the LLM may not always be enough. Note that it is common to have multiple rounds of discussion even between the project maintainers before a pull request is created. Hence, a flexible interface and human involvement criteria between human developers and LLM agents are desired. When an LLM agent gets stuck in any stage of the workflow, the human developers can be asked to provide additional insights on the issue which may steer the LLM agent to the correct reasoning direction and eventually resolve more challenging tasks.
@@ 8. Threats to Validity
Despite the high efficacy AutoCodeRover achieved in SWE-bench, we report a few potential threats to our approach and experiments and discuss how we addressed them. First, LLM is known for its randomness to generate different results in different runs, which may threaten the validity of AutoCodeRover’s performance. We address this by repeating our main evaluation experiment three times and releasing a replication package for practitioners. Second, the operating system environment may affect the evaluation result for the project subjects in SWE-bench (e.g., matplotlib). We share the evaluation environment in Section 5 and provide a docker environment for usage by the community. Third, we consider a task instance in SWE-bench as resolved if the generated patch passes all the test cases added in the pull request for the issue. However, this patch may be overfitting. We addressed this threat by manually verifying whether the patches are semantically equivalent to developers’ patches among the authors. Finally, our result of Swe-agent-rep is lower than reported by SWE-bench team. This might be due to the different OS environments, different LLM, and parameter setups. To address this, we replicated Swe-agent-rep with the default setting in the provided scripts, and we used the same LLM (gpt-4-0125-preview) on Swe-agent-rep as AutoCodeRover. We also ran experiments in the docker provided by SWE-bench team (Table 3); this docker was publicly released on 6 April 2024.
@@ 9. Perspectives
AI-based software engineering is currently a topic of study among researchers, innovators, and entrepreneurs. Part of the trigger for this interest has been provided by efforts like GitHub Copilot (gco, 2022) in the last few years. These efforts show significant promise in the use of Large Language Models (LLMs) to automatically generate code. At the same time, code generated by LLMs may be incorrect (Fan et al., 2023) or vulnerable (Pearce et al., 2023). Thus, we need autonomous processes which allow for code improvements, such as bug fixes and feature additions.
In this paper, we suggest a LLM-based solution AutoCodeRover for autonomous code improvement. The key distinguishing feature of AutoCodeRover is its conscious attempt to inject a software engineering outlook by integrating (a) use of program representations such as ASTs instead of files, (b) iterative code search by exploiting program structure and (c) use of test-based fault localization when tests can be constructed. AutoCodeRover shows significant efficacy in terms of solving real-life Github issues. AutoCodeRover shows that relying on the GitHub issue description to guide code search for patches / modifications can be misleading. It shows the need for use of program representation and structure to enable autonomous code modifications.
Today, the code LLMs cannot produce safe and secure code which can be trusted enough to be integrated into real software projects. There is thus a need to autonomously improve code (both automatically generated and manually written), for which LLMs can play a role. Future work needs to focus further on the appropriate points when tools like AutoCodeRover may converse with the human programmer. Developers may need to shift to playing different roles at the same time in future software industry - vetting different conversations with LLM-based tools like AutoCodeRover to enable a variety of software engineering activities. This would contrast with today’s software industry where a person has specific roles like programmer/tester/architect/requirements-engineer. Thus, apart from full-stack engineers, we may see more full-lifecycle software engineers in future, who are comfortable to work with the entire lifecycle of (the components of) a software system. Furthermore, the focus of LLM oriented workflows is on scale today. This focus may shift to engendering of trust if LLM oriented autonomous software engineering becomes commonplace. Future software engineers may focus more on greater trust, instead of larger scale.
@@
We share full public access to (1) AutoCodeRover’s implementation, (2) all patches generated during our experiment and conversation history with LLM, (3) replication scripts for our experiments, (4) replication scripts and logs for Swe-agent-rep at the following website: https://github.com/nus-apr/auto-code-rover
@@ Acknowledgements.
This work was partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant ”Automated Program Repair”, MOE-MOET32021-0001.

### v3

@@ 6.2. RQ2: Effect of SBFL
Bug reproduction (Jin and Orso, 2012; Böhme et al., 2017) is a well-studied topic in the software engineering community. It aims to automatically construct input tests to reproduce bugs that are described in bug reports. Subsequently, those bug reproduction tests can be used by program analysis techniques to localize the root cause of bugs and generate patches for APR tools (Nguyen et al., 2013).
In this research question, we investigate whether program analysis techniques with reproducible tests can benefit the workflow of AutoCodeRover, and we use spectrum-based fault localization (SBFL) as an example. Different from RQ1, here we construct a common scenario in program repair that AutoCodeRover has access to the complete test-suite of the target task instance. We use the developer-written test cases for each task instance (provided in SWE-bench lite) as the test-suite. To understand the effect of SBFL, we additionally provide SBFL results (top-5 suspicious methods) to AutoCodeRover at the beginning of the context retrieval stage, then use the test-suite for patch validation during the patch generation retry-loop. We denote this setting as ACR-sbfl in Table 2. The patch validation is added as follows: when a patch is generated by the LLM agent, the test-suite is executed on the patched program; if the patch fails to pass the complete test-suite, AutoCodeRover re-invokes the patch generation agent to write a new patch. This validation loop is configured to run three times.
@@ Results.
Table 2 shows that, with the additional information provided by SBFL, the number of resolved tasks increased from 57 to 66 (i.e. 19% to 22% resolved rate on SWE-bench lite). Moreover, when comparing with task instances resolved in AutoCodeRover on SWE-bench lite, ACR-sbfl still uniquely resolves 7 task instances that are not resolved in any of the other runs.
@@
We present a case study on one of the tasks uniquely resolved by ACR-sbfl. Figure 9 shows the issue description of django-13964555https://code.djangoproject.com/ticket/32332 and how AutoCodeRover attempts to retrieve code context with and without SBFL. This issue reported a bug when saving django models to a database. A simplified version of the issue description is shown in the first part of Figure 9, in which important parts are highlighted. In this issue, some “hints” (highlighted in red) mentioned are actually distracting factors for a context retrieval agent. For example, the Product and Order classes describe how the bug can be reproduced, and are not classes that cause the bug. Retrieving code context from these two classes (as shown in Figure 9 part 2 “WITHOUT SBFL”) does not yield useful results for forming the context and resolving the issue.
@@ 6.3. RQ3: Challenges on real-life tasks
In this research question, we analyze the task instances in SWE-bench lite that AutoCodeRover failed to resolve (based on the result of ACR-all in Table 2, without applying SBFL) and provide a taxonomy of the issue characteristics to highlight the practical challenges in achieving fully automated software improvement. Our taxonomy consists of challenges in the fault localization stage and patch generation stage. Specifically, for each task, we analyze the best run in the three repetitions, and classify each of the 300 tasks into one of the following:
@@
Figure 10 shows the distribution of the 300 tasks in SWE-bench lite. AutoCodeRover resolves 26.0% of the issues (“Success”), as mentioned in Section 6.1. The fail-to-resolve cases are included in the remaining four categories. In 29.3% of the tasks, AutoCodeRover correctly decided on all patch locations (at the method-level), but did not produce a correct patch (“Wrong patch”). More fine-grained intra-procedural analysis and specification inference techniques can play a significant role in improving these cases, by providing the patch generation agent with more method-level repair guidance. In the other three categories, the fault localization could not pinpoint all the locations to be modified. In 20.0% of the tasks, a patch is generated in the correct file, but at wrong methods / classes in the file (“Wrong location in correct file”). In some of these runs, the developer patch modifies multiple methods, but the generated patch did not modify all of them. In the other categories, a patch could not be generated at the correct file - in 18.0% of the tasks a patch is generated in wrong files, and in 6.7% of the tasks there is no applicable patch (“Wrong file” and “No patch”). We manually inspected some tasks in these two categories, and observed that their issue description mentions few methods / classes / files in the codebase. Instead, some of them contain short examples to reproduce the issue. For these tasks, one possibility is to generate a comprehensive test-suite based on the issue description, and then use execution information of the test-suite (e.g. SBFL) to reveal suspicious program locations. On the other hand, some other tasks do not contain reproducible examples and only consist of natural language descriptions. For these tasks some human involvement might be helpful. The developers could focus on these tasks.
7. Discussion on Experiments and Improvements
In this section, we discuss our position on the experiment results and a few possible directions for future improvements.
Position on Experiment Results
Building autonomous large language model agent systems for software engineering tasks is one of the fastest-growing research fields now. There were more than 17 attempts from both academia and industries on SWE-bench since April 2024. We refer the practitioners to the SWE-bench Leaderboard (Jimenez et al., 2024a) (refer Figure 11) which maintains the latest research effort of various recent agents. The most recent resolve rate on full SWE-bench leaderboard has reached 19.27% by Factory Code Droid as of 12th July 2024. In the meantime, AutoCodeRover is also rapidly progressing towards better performance, the most recent update of AutoCodeRover (18.83% on SWE-bench on 12th July 2024) can also be found at (aut, 2024). However, despite the satisfying results of AutoCodeRover, it is useful to interpret the results in the right spirit. Apart from showing good efficacy with low cost, AutoCodeRover is significant because of the way it tries to generate program modifications. Thus, we highlight the importance of gleaning program specifications to guide the patching process by agents as a general guideline for future research.
Figure 11. Snapshot of full SWE-bench leaderboard on 12th July 2024
@@ 7. Discussion on Experiments and Improvements
In Section 4, we described two scenarios that AutoCodeRover can work on, (1) issue description only, (2) aided by SBFL when the test suite is available. Although the second scenario is not always practical in real-life software development, the issue description sometimes contains a concrete reproduction script from the user that reports the bug. In the future, it is possible to design an LLM agent specifically to generate a bug reproduction test based on the GitHub issue description. The bug reproduction test can then be used to validate the correctness of AutoCodeRover’s generated patch and possibly enable a regeneration process if a patch fails to pass the reproduction test.
@@
During context retrieval, AutoCodeRover effectively navigates through the codebase by visiting code entities such as classes and methods. The idea of utilizing code-specific structure for context retrieval can be taken further by considering artifacts from program semantics. For example, from an initial set of methods identified in the issue description, a static call graph analysis (Ryder, 1979; Salis et al., 2021) can be used to collect additional relevant methods when testcases are absent. There are also potential in integrating a language server (Gunasinghe and Marcus, 2021) for codebase navigation, so that the context retrieval agent can perform more code-specific actions such as “jump-to-definition” from a method invocation. Finally, one can use forward data dependence analysis (Jin et al., 2024) from the methods in the issue description to search for other relevant methods.
@@
Currently, AutoCodeRover left all decision-making processes to its foundation LLM — e.g. deciding context retrieving locations and whether to terminate the retrieval stage. However, leaving it to the LLM may not always be enough. Note that it is common to have multiple rounds of discussion even between the project maintainers before a pull request is created. Hence, a flexible interface and human involvement criteria between human developers and LLM agents are desired.
@@ 8. Threats to Validity
Despite the high efficacy AutoCodeRover achieved in SWE-bench, we report a few potential threats to our approach and experiments and discuss how we addressed them. First, LLM is known for its randomness to generate different results in different runs, which may threaten the validity of AutoCodeRover’s performance. We address this by repeating our main evaluation experiment three times and releasing a replication package for practitioners. Second, we consider a task instance in SWE-bench as resolved if the generated patch passes all the test cases added in the pull request for the issue. However, this patch may be overfitting. We addressed this threat by manually verifying whether the patches are semantically equivalent to developers’ patches among the authors.
@@ 9. Perspectives
AI-based software engineering is currently a topic of study among researchers, innovators, and entrepreneurs. Part of the trigger for this interest has been provided by efforts like GitHub Copilot (gco, 2022) in the last few years. These efforts show significant promise in the use of Large Language Models (LLMs) to automatically generate code. At the same time, code generated by LLMs may be incorrect (Fan et al., 2023) or vulnerable (Pearce et al., 2023b). Thus, we need autonomous processes which allow for code improvements, such as bug fixes and feature additions.
In this paper, we suggest a LLM-based solution AutoCodeRover for autonomous code improvement. The key distinguishing feature of AutoCodeRover is its conscious attempt to inject a software engineering outlook by integrating (a) use of program representations such as ASTs instead of files, (b) iterative code search by exploiting program structure and (c) use of test-based fault localization when tests can be constructed. AutoCodeRover shows significant efficacy in terms of solving real-life GitHub issues. AutoCodeRover shows that relying only on the GitHub issue description to guide code search for patches / modifications can be misleading. From the point of view of program repair, the toughest challenge lies in inferring the intent of the developer or the specification. The outlook of AutoCodeRover is that the developer intent is gleaned from the project structure for automated program modifications.
Today, the code LLMs cannot produce safe and secure code which can be trusted enough to be integrated into real software projects. There is thus a need to autonomously improve code (both automatically generated and manually written), for which LLMs can play a role. Future work needs to focus further on the appropriate points when tools like AutoCodeRover may converse with the human programmer. Developers may need to shift to playing different roles at the same time in future software industry - vetting different conversations with LLM-based tools like AutoCodeRover to enable a variety of software engineering activities. This would contrast with today’s software industry where a person has specific roles like programmer/tester/architect/requirements-engineer. Thus, apart from full-stack engineers, we may see more full-lifecycle software engineers in the future, who are comfortable working with the entire lifecycle of (the components of) a software system. Furthermore, the focus of LLM-oriented workflows is on scale today. This focus may shift to the engendering of trust if LLM-oriented autonomous software engineering becomes commonplace. Future software engineers may focus more on greater trust, instead of larger scale.
@@
We share full public access to (1) AutoCodeRover’s implementation, (2) all patches generated during our experiment and conversation history with LLM, (3) replication scripts for our experiments at the following website: https://github.com/nus-apr/auto-code-rover.
@@ Acknowledgements.
We thank the anonymous reviewers for their suggestions. This work was partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant ”Automated Program Repair”, MOE-MOET32021-0001.